# Stochastic Monkeys at Play: Random Augmentations Cheaply Break LLM Safety Alignment

### v1

@@ Abstract
Safety alignment of Large Language Models (LLMs) has recently become a critical objective of model developers. In response, a growing body of work has been investigating how safety alignment can be bypassed through various jailbreaking methods, such as adversarial attacks. However, these jailbreak methods can be rather costly or involve a non-trivial amount of creativity and effort, introducing the assumption that malicious users are high-resource or sophisticated. In this paper, we study how simple random augmentations to the input prompt affect safety alignment effectiveness in state-of-the-art LLMs, such as Llama 3 and Qwen 2. We perform an in-depth evaluation of 17 different models and investigate the intersection of safety under random augmentations with multiple dimensions: augmentation type, model size, quantization, fine-tuning-based defenses, and decoding strategies (e.g., sampling temperature). We show that low-resource and unsophisticated attackers, i.e. stochastic monkeys, can significantly improve their chances of bypassing alignment with just 25 random augmentations per prompt.
@@ 1 Introduction
    We also observe that: 1 Character-level augmentations tend to be much more effective than string insertion augmentations for increasing success rate, 2 Larger models tend to be safer, 3 More aggressive weight quantization tends to be less safe, 4 Adversarial training can generalize to random augmentations, but its effect can be circumvented by decreasing augmentation intensity, and 5 Even when altering the sampling temperature, random augmentations still provide further improvements to the success rate. We also employ a human study on a sample of 1220 data points from our experiments to calibrate our evaluation metric for controlling the estimated false positive and false negative rates.
@@ 2.3 Evaluation Metric
Given we use a learned classifier for $c$, simply checking if any (i.e., $\gamma=0$) augmentation succeeds can have a high false positive rate (a false positive occurs when $s_{k,\gamma}(\textbf{x},\textbf{y}_{1},\ldots,\textbf{y}_{k})$ evaluates to 1 when in fact none of the $k$ outputs are harmful). A non-zero $\gamma$ can therefore be used to help reduce the false positive rate. However, applying too high of a threshold may result in a high false negative rate (a false negative occurs when $s_{k,\gamma}(\textbf{x},\textbf{y}_{1},\ldots,\textbf{y}_{k})$ evaluates to 0 when in fact at least one of the $k$ outputs are harmful). Thus, $\gamma$ should be carefully chosen so as to balance the false positive and false negative rates. See Appendix B.1 for more details.
@@ 3 Experimental Setup
For computing $(k,\gamma)$-success rates, we set $k=25$ to reduce the runtime of experiments and since we find this value to be sufficient for significantly affecting the success rate. Since the $(k,\gamma)$-success false positive and false negative rates may vary significantly for each augmentation set $\mathcal{A}$, we use separate thresholds $\gamma^{*}_{\mathcal{A}}$ that balances empirical estimates of the false positive and false negative rates for each $\mathcal{A}$, and employ a human study to obtain these empirical estimates (see Appendix B.2 for more details on selecting $\gamma^{*}_{\mathcal{A}}$). As different model developers may impose different objectives for safety alignment, we seek a dataset $\mathcal{D}$ that covers a broad range of harmful behavior classes. Hence, we use the SORRY-Bench dataset (Xie et al., 2024), which was constructed by sampling 450 prompts from across 10 datasets in a class-balanced manner over 45 classes. For the safety judge $c$, we use the fine-tuned Mistral-based safety judge that accompanies the SORRY-Bench dataset.
@@ 3.1 Augmentations
For the RQ1 experiments, we examine the success rate gain $\hat{r}_{25,\gamma^{*}_{\mathcal{A}}}(\mathcal{A},f,g)-\hat{r}_{25,\gamma^{*}_{\mathcal{A}}}(\mathcal{A}_{I},f,g)$ of using an augmentation set $\mathcal{A}$ over using no augmentations. Note that since adversarial attacks typically assume white-box access to the model, and since random augmentations are an entirely black-box operation, we do not compare against adversarial attack performance.
@@ 3.1.2 Augmentation Strength
For string insertion augmentations, the notion of augmentation “strength” refers to the length of the inserted string, whereas for character-level augmentations, “strength” refers to the amount of character positions that are augmented. We consider two ways to control the strength of an augmentation: 1. The strength is fixed for each prompt, and 2. The strength is proportional to the length of each prompt. Since $\mathcal{D}$ may contain a wide range of prompt lengths, fixing the strength may result in augmentations that are too aggressive for short prompts (which may change their semantic meaning) or too subtle for long prompts (which may lead to low success rate gains), in particular for character-level augmentations. Therefore, we focus on proportional augmentation strength, as governed by a proportion parameter $p$. For instance, with $p=0.1$ and an original prompt length of 200 characters, the inserted string length for string insertion augmentations and the amount of augmented character positions for character-level augmentations would be 20 characters. (The number of characters is always rounded down to the nearest integer.) For our experiments, we set $p=0.05$, which we find to be sufficient for obtaining non-trivial success rate gains while ensuring the augmentations are not too aggressive for shorter prompts (see the examples in Table 1, which use $p=0.05$).
@@ 3.2 Models
We consider the following models across 8 different model families: Llama 2 (Llama 2 7B Chat, Llama 2 13B Chat) (Touvron et al., 2023), Llama 3 (Llama 3 8B Instruct) (Dubey et al., 2024), Llama 3.1 (Llama 3.1 8B Instruct), Mistral (Mistral 7B Instruct v0.2), Phi 3 (Phi 3 Mini 4K Instruct, Phi 3 Small 8K Instruct, Phi 3 Medium 4K Instruct), Qwen 2 (Qwen 2 0.5B, Qwen 2 1.5B, Qwen 2 7B), Vicuna (Vicuna 7B v1.5, Vicuna 13B v1.5) and Zephyr (Zephyr 7B Beta). Among these, only the Llama, Phi and Qwen families have undergone explicit safety alignment. The remaining families are included to see if any interesting patterns can be observed for unaligned models. For instance, although Mistral was not explicitly aligned, it can sometimes exhibit refusal behavior for some harmful prompts, so it would still be interesting to see how this behavior can be affected by random augmentations.
@@ 3.3 Decoding Strategies
By default, all our experiments are conducted using greedy decoding to isolate the randomness effects of using multiple random augmentations. However, for the RQ3 experiments, for each augmentation set $\mathcal{A}$ we examine the success rate gain $\hat{r}_{25,\gamma^{*}_{\mathcal{A}}}(\mathcal{A},f,g)-\hat{r}_{25,\gamma^{*}_{\mathcal{A}}}(\mathcal{A}_{I},f,g)$ for sampling-based generation algorithms $g$. Specifically, we consider temperature sampling with various temperatures $\tau$ for $g$. We consider two values for $\tau$: 0.7 (since this is a value in the range of common temperature parameters between 0.6 and 0.9), and 1.0 (to explore the largest possible temperature value). Note that greedy decoding is a special case where $\tau=0.0$.
@@ 4 Experimental Results
In this section, we plot the results for each of our experiments and discuss our observations. Raw data values (including results using a fixed $\gamma=0$ for all augmentations) broken down by augmentation type are reported in Appendix C.
@@ 4.1 RQ1: Varying Augmentation Type
Figure 2: Average $\left(25,\gamma^{*}_{\mathcal{A}}\right)$-success rate gains of different kinds of augmentations over using no augmentations for various sampling temperatures $\tau$. RQ1 results are where $\tau=0.0$, and the rest are RQ3 results.
@@ 4.2.2 Quantization
Figure 4 reports the quantization experiment results for RQ2. For W8A8, most success rate changes are small, with all deviations being within 5%. Among all models, Qwen 2 7B has the greatest tendency towards becoming less safe. In Figure 12 in Appendix D, we show an example where the original Qwen 2 model fails under the random suffix augmentation while the W8A8 model succeeds even when the random suffixes used are the exact same for both models. Moving over to the W4A16 results, we see that the Llama 3, Mistral, Phi and Vicuna models become noticeably less safe. However, other models such as Llama 2, Llama 3.1 and Zephyr barely change, similar to their W8A8 counterparts. Even more curiously however, we see that Qwen 2 7B seemingly becomes more safe. However, upon further inspection, we realize that this may be a result of poorer model response quality in general; see Figure 13 in Appendix D for examples. Overall, while quantization can have some significant influence on success rate with more aggressive weight quantization tending to reduce safety, these effects are not consistent across models. As with the results for the model size experiment, this suggests that there may be other underlying factor(s) that determine how quantization affects safety under random augmentations.
@@ 4.2.3 Fine-Tuning-Based Defense
To provide a fairer assessment of the improvements afforded by the adversarial training, we also examine fixed-length random suffixes in Figure 6. Note that character length 25 has an average token length of ${\sim}22$, which is close to the fixed length of 20 tokens R2D2 used. The blue curve shows that as the length increasess, the success rate continues to decrease, even somewhat past length 25. This suggests that length generalization is unidirectional: protection is afforded to lengths beyond the adversarial training length, but starts to disappear for shorter suffixes. This heavily contrasts with adversarial training in the image classification domain, where protections typical only hold with image perturbations less extreme than what was trained against (Gowal et al., 2018).
@@ 4.3 RQ3: Varying the Generation Configuration
Figure 2 reports the experiment results for RQ3 (denoted by $\tau=0.7$ and $\tau=1.0$). First, we remark that increasing temperature without any augmentations already increases the success rate (see Table 3); this is in line with the findings of Huang et al. (2023) that showed altering temperature alone can be a successful attack. Next, we observe, that across all models and for both values of $\tau$, using random augmentations further improves the success rate on top of success gains due to output sampling. This supplements the findings of Huang et al. (2023) which only explored success rate improvements for prompts without any input augmentations. Our results show that two sources of randomness, namely output sampling and input augmentations, can work together to provide even greater attack effectiveness.
@@ 5 Conclusion
This paper demonstrates that simple random augmentations are a cheap yet effective approach to bypassing the safety alignment of state-of-the-art LLMs. However, although we identify patterns in how dimensions such as model size and quantization affect safety under random augmentations, we also discover counterexamples to these patterns. As such, the full story may be much more complex and involve additional underlying factors such as training data, optimization, etc. In the future, we will investigate whether these more complex dimensions can explain the counterexamples we observed. Moreover, we will go beyond evaluation to identify ways that models can become more robust against random augmentations.
@@ Appendix A Related Work
A.1 Random Augmentations and Robustness
@@
A.2 Random Augmentations in Adversarial Attacks
Random augmentations have also been utilized within adversarial attack algorithms. For instance, Li et al. (2018) introduced the TextBugger attack framework, which adversarially applies random augmentations (e.g., character-level augmentations such as inserting, swapping, or deleting characters and word-level augmentations such as word substitution) to fool models on sentiment analysis, question answering and machine translation tasks. Their method computes a gradient to estimate word importance, and then uses this estimate to apply random augmentations at specific locations based on the importance estimation. Additionally, Morris et al. (2020) introduced a comprehensive framework for generating adversarial examples to attack NLP models such as BERT, utilizing the word-level augmentations from the Easy Data Augmentation method  (Wei & Zou, 2019) (i.e., synonym replacement, insertion, swapping, and deletion). The adversarial examples are also used to perform adversarial training to improve model robustness and generalization.
A.3 Random Augmentations for Defense
SmoothLLM  (Robey et al., 2023) was introduced as a system-level defense for mitigating jailbreak effectiveness. Their key observation is that successful jailbreaks are extremely brittle to random augmentations; i.e., many of the successful jailbreaks won’t succeed after augmentation. In contrast, our work is based on the observation that the original prompt itself is also brittle, but in the opposite direction: given a prompt that doesn’t succeed, one can effectively find an augmented prompt that does succeed. Moreover, their attack success evaluation is only based on a single chosen output per prompt, effectively discarding the other $k-1$ outputs. In contrast, since our threat model is built around the attacker making $k$ independent attempts per prompt, our attack success evaluation accounts for all of the $k$ outputs per prompt.

### v2

@@ Abstract
Safety alignment of Large Language Models (LLMs) has recently become a critical objective of model developers. In response, a growing body of work has been investigating how safety alignment can be bypassed through various jailbreaking methods, such as adversarial attacks. However, these jailbreak methods can be rather costly or involve a non-trivial amount of creativity and effort, introducing the assumption that malicious users are high-resource or sophisticated. In this paper, we study how simple random augmentations to the input prompt affect safety alignment effectiveness in state-of-the-art LLMs, such as Llama 3 and Qwen 2. We perform an in-depth evaluation of 17 different models and investigate the intersection of safety under random augmentations with multiple dimensions: augmentation type, model size, quantization, fine-tuning-based defenses, and decoding strategies (e.g., sampling temperature). We show that low-resource and unsophisticated attackers, i.e. stochastic monkeys, can significantly improve their chances of bypassing alignment with just 25 random augmentations per prompt. Source code and data: https://github.com/uiuc-focal-lab/stochastic-monkeys/.
@@ 1 Introduction
    We also observe that: 1 Character-level augmentations tend to be much more effective than string insertion augmentations for increasing success rate, 2 Larger models tend to be safer, 3 More aggressive weight quantization tends to be less safe, 4 Adversarial training can generalize to random augmentations, but its effect can be circumvented by decreasing augmentation intensity, and 5 Even when altering the sampling temperature, random augmentations can sometimes provide further success rate improvement. We also employ a human study on a sample of 1220 data points from our experiments to calibrate our evaluation metric for controlling the estimated false positive and false negative rates.
@@ 2.3 Evaluation Metric
Given we use a learned classifier for $c$, simply checking if any (i.e., $\gamma=0$) augmentation succeeds can have a high false positive rate (a false positive occurs when $s_{k,\gamma}(\textbf{x},\textbf{y}_{1},\ldots,\textbf{y}_{k})$ evaluates to 1 when in fact none of the $k$ outputs are harmful). A non-zero $\gamma$ can therefore be used to help reduce the false positive rate. However, applying too high of a threshold may result in a high false negative rate (a false negative occurs when $s_{k,\gamma}(\textbf{x},\textbf{y}_{1},\ldots,\textbf{y}_{k})$ evaluates to 0 when in fact at least one of the $k$ outputs are harmful). Thus, $\gamma$ should be carefully chosen so as to balance the false positive and false negative rates. See Appendix C.1 for more details.
@@ 3 Experimental Setup
For computing $(k,\gamma)$-success rates, we set $k=25$ to reduce the runtime of experiments and since we find this value to be sufficient for significantly affecting the success rate. Since the $(k,\gamma)$-success false positive and false negative rates may vary significantly for each augmentation set $\mathcal{A}$, we use separate thresholds $\gamma^{*}_{\mathcal{A}}$ that balances empirical estimates of the false positive and false negative rates for each $\mathcal{A}$, and employ a human study to obtain these empirical estimates (see Appendix C.2 for more details on selecting $\gamma^{*}_{\mathcal{A}}$). As different model developers may impose different objectives for safety alignment, we seek a dataset $\mathcal{D}$ that covers a broad range of harmful behavior classes. Hence, we use the SORRY-Bench dataset (Xie et al., 2024), which was constructed by sampling 450 prompts from across 10 datasets in a class-balanced manner over 45 classes. For the safety judge $c$, we use the fine-tuned Mistral-based safety judge that accompanies the SORRY-Bench dataset.
@@ 3.1 Augmentations
For the RQ1 experiments, we examine the success rate gain $\hat{r}_{25,\gamma^{*}_{\mathcal{A}}}(\mathcal{A},f,g)-\hat{r}_{25,\gamma^{*}_{\mathcal{A}}}(\mathcal{A}_{I},f,g)$ of using an augmentation set $\mathcal{A}$ over using no augmentations. See Appendix D.4 for a comparison of RQ1 results against PAIR (Chao et al., 2023) and DeepInception (Li et al., 2023) attacks.
@@ 3.1.2 Augmentation Strength
For string insertion augmentations, the notion of augmentation “strength” refers to the length of the inserted string, whereas for character-level augmentations, “strength” refers to the amount of character positions that are augmented. We consider two ways to control the strength of an augmentation: 1. The strength is fixed for each prompt, and 2. The strength is proportional to the length of each prompt. Since $\mathcal{D}$ may contain a wide range of prompt lengths, fixing the strength may result in augmentations that are too aggressive for short prompts (which may change their semantic meaning) or too subtle for long prompts (which may lead to low success rate gains), in particular for character-level augmentations. Therefore, we focus on proportional augmentation strength, as governed by a proportion parameter $p$. For instance, with $p=0.1$ and an original prompt length of 200 characters, the inserted string length for string insertion augmentations and the amount of augmented character positions for character-level augmentations would be 20 characters. (The number of characters is always rounded down to the nearest integer.) For our experiments, we set $p=0.05$, which we find to be sufficient for obtaining non-trivial success rate gains while ensuring the augmentations are not too aggressive for shorter prompts (see Table 1). See Appendix D.3 for an ablation study on $p$.
@@ 3.2 Models
We consider the following models across 8 different model families: Llama 2 (Llama 2 7B Chat, Llama 2 13B Chat) (Touvron et al., 2023), Llama 3 (Llama 3 8B Instruct) (Dubey et al., 2024), Llama 3.1 (Llama 3.1 8B Instruct), Mistral (Mistral 7B Instruct v0.2), Phi 3 (Phi 3 Mini 4K Instruct, Phi 3 Small 8K Instruct, Phi 3 Medium 4K Instruct), Qwen 2 (Qwen 2 0.5B, Qwen 2 1.5B, Qwen 2 7B), Vicuna (Vicuna 7B v1.5, Vicuna 13B v1.5) and Zephyr (Zephyr 7B Beta). In Appendix D.1, we also evaluate GPT-4o OpenAI (2024). Among these, only the Llama, Phi and Qwen families have undergone explicit safety alignment. The remaining families are included to see if any interesting patterns can be observed for unaligned models. For instance, Mistral can sometimes exhibit refusal behavior for harmful prompts, so it would still be interesting to see how this is be affected by random augmentations. By default, we leave the system prompt empty for all models; see Appendix D.2 for an experiment with safety-encouraging system prompts.
@@ 3.3 Decoding Strategies
By default, all our experiments are conducted using greedy decoding to isolate the randomness effects of using multiple random augmentations. However, for the RQ3 experiments, for each augmentation set $\mathcal{A}$ we examine the success rate gain $\hat{r}_{25,\gamma^{*}_{\mathcal{A}}}(\mathcal{A},f,g)-\hat{r}_{25,\gamma^{*}_{\mathcal{A}}}(\mathcal{A}_{I},f,g)$ for sampling-based generation algorithms $g$. Specifically, we consider temperature sampling with various temperatures $\tau$ for $g$. We consider two values for $\tau$: 0.7 (since this is a value in the range of common temperature parameters between 0.6 and 0.9), and 1.0 (to explore the largest possible temperature value). We set the maximum generated tokens to be 1024.
@@ 4 Experimental Results
In this section, we plot the results for each of our experiments and discuss our observations. Raw data values (including results using a fixed $\gamma=0$ for all augmentations) broken down by augmentation type are reported in Appendix D. Examples of successful attacks can be found in Appendix D.5.
@@ 4.1 RQ1: Varying Augmentation Type
Figure 2: Average $\left(25,\gamma^{*}_{\mathcal{A}}\right)$-success rate gains of different kinds of augmentations over using no augmentations, using greedy decoding for generation.
@@ 4.2.2 Quantization
Figure 4 reports the quantization experiment results for RQ2. For W8A8, most success rate changes are small, with all deviations being within 5%. Among all models, Qwen 2 7B has the greatest tendency towards becoming less safe. In Figure 18 in Appendix D.5, we show an example where the original Qwen 2 model fails under the random suffix augmentation while the W8A8 model succeeds even when the random suffixes used are the exact same for both models. Moving over to the W4A16 results, we see that the Llama 3, Llama 3.1, Mistral, Phi and Vicuna models become noticeably less safe. However, Llama 2 and Zephyr barely change, similar to their W8A8 counterparts. Even more curiously however, we see that Qwen 2 7B seemingly becomes more safe. However, upon further inspection, we realize that this may be a result of poorer model response quality in general; see Figure 19 in Appendix D.5 for examples. Overall, while quantization can have some significant influence on success rate with more aggressive weight quantization tending to reduce safety, these effects are not consistent across models. As with the results for the model size experiment, this suggests that there may be other underlying factor(s) that determine how quantization affects safety under random augmentations.
@@ 4.2.3 Fine-Tuning-Based Defense
To provide a fairer assessment of the improvements afforded by the adversarial training, we also examine fixed-length random suffixes in Figure 6. Note that character length 25 has an average token length of ${\sim}22$, which is close to the fixed length of 20 tokens R2D2 used. The blue curve shows that as the length increases, the success rate continues to decrease, even somewhat past length 25. This suggests that length generalization is unidirectional: protection is afforded to lengths beyond the adversarial training length, but starts to disappear for shorter suffixes. This heavily contrasts with adversarial training in the image classification domain, where protections typical only hold with image perturbations less extreme than what was trained against (Gowal et al., 2018).
@@ 4.3 RQ3: Varying the Generation Configuration
Figure 12 and Table 7 in Appendix D report the experiment results for RQ3. First, we remark that increasing temperature without any augmentations already increases the success rate; this is in line with the findings of Huang et al. (2023) that showed altering temperature alone can be a successful attack. Next, we observe that applying random augmentations on top of output sampling overall tends to hurt the success rate. However, from Table 7, we see that for Llama 2, Llama 3 and Phi 3, character deletion further improves the success rate. This shows that two sources of randomness, namely output sampling and random augmentations, can sometimes work together to provide even greater attack effectiveness.
4.4 Discussion
In summary, we provide a ranking for how influential each dimension is on safety: 1. Fine-tuning-based defenses; e.g., Mistral 7B with RR experiences a 55.9% improvement in safety on average (see Table 10), 2. Model size; e.g., Qwen 2 0.5B drops 23.2% in safety from 1.5B on average (see Table 8), 3. Quantization; while W8A8 maintains safety, W4A16 tends to reduce it (e.g., with Llama 3 dropping 10.5%), and 4. Output sampling, which only rarely decreases safety (and tends to improve it). Please see Appendix B for discussion on the practical implications of random augmentation attacks.
@@ 5 Conclusion
This paper demonstrates that simple random augmentations are a cheap yet effective approach to bypassing the safety alignment of state-of-the-art LLMs. Our work aims to add a broader characterization of this specific vulnerability to the ongoing discussion of LLM safety. As such, through exploring a diverse set of models and random augmentations, we identify general trends in how dimensions such as model size and quantization affect safety under random augmentations. Future work can investigate how more complex dimensions such as training data and optimization interact with LLM safety under random augmentations, as well as dive deeper into explaining why LLM safety can be so brittle to small character-level augmentations.
@@ Appendix A Related Work
A.1 Simple Techniques for Bypassing Safety Alignment
A growing number of simple techniques for bypassing safety alignment have recently been proposed. These methods are simpler in comparison to adversarial attacks such as GCG (Zou et al., 2023), but may also involve threat models that have different assumptions about the attacker. Huang et al. (2023) showed that searching over different decoding configuration can yield model responses that bypass safety alignment; the attacker only needs to have the ability to alter the generation configuration, and therefore this technique can be more readily applied to closed-source models (e.g., through API access). Andriushchenko & Flammarion (2024) showed that rephrasing a prompt into the past tense can also successfully jailbreak LLMs. This involves even fewer assumptions about the attacker, and the conversion to past tense can either be performed manually with relative ease (or automated with another LLM for mass evaluation). Vega et al. (2023) showed that the safety alignment of open-source models can be easily bypassed by simply prefilling the assistant response with a compliant string in what are now known as prefilling attacks. More generally, this assumes that the attacker has prefilling access, which is offered by some closed-source models such as Claude through their API Andriushchenko et al. (2024). In contrast to these works, the random augmentations we explore in our work involves very few assumptions about the attacker (i.e., only requiring black-box access), and can be easily applied to prompts programmatically (i.e., not requiring any manual effort or auxiliary LLMs).
A.2 Random Augmentations and Robustness
@@
A.3 Random Augmentations in Adversarial Attacks
Techniques that use random augmentations for attack purposes have largely focused on using the random augmentations as part of a larger adversarial attack algorithm, rather than simply using the random augmentations as an attack in itself. For instance, Li et al. (2018) introduced the TextBugger attack framework, which adversarially applies random augmentations (e.g., character-level augmentations such as inserting, swapping, or deleting characters and word-level augmentations such as word substitution) to fool models on sentiment analysis, question answering and machine translation tasks. Their method computes a gradient to estimate word importance, and then uses this estimate to apply random augmentations at specific locations based on the importance estimation. Additionally, Morris et al. (2020) introduced a comprehensive framework for generating adversarial examples to attack NLP models such as BERT, utilizing the word-level augmentations from the Easy Data Augmentation method  (Wei & Zou, 2019) (i.e., synonym replacement, insertion, swapping, and deletion). The adversarial examples are also used to perform adversarial training to improve model robustness and generalization.
A.4 Random Augmentations for Defense
Of the comparatively fewer works that investigate random augmentations in the context of generative language model safety, most focus on applying augmentations for defense purposes. For example, SmoothLLM  (Robey et al., 2023) was introduced as a system-level defense for mitigating jailbreak effectiveness. Their key observation is that successful jailbreaks are extremely brittle to random augmentations; i.e., many of the successful jailbreaks won’t succeed after augmentation. In contrast, our work is based on the observation that the original prompt itself is also brittle, but in the opposite direction: given a prompt that doesn’t succeed, one can effectively find an augmented prompt that does succeed. Moreover, their attack success evaluation is only based on a single chosen output per prompt, effectively discarding the other $k-1$ outputs. In contrast, since our threat model is built around the attacker making $k$ independent attempts per prompt, our attack success evaluation accounts for all of the $k$ outputs per prompt.